What volume of CCl4 having density 1.5g/cc contains 1 * 10 to the power 25 chlorine atoms

Grade:

2 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

Dear student,

number of chlorine atoms in ccl4 = 4

number of moles of ccl4 = 10^25/(4*6.623*10^23) = 3.775 moles

weight of 3.775 moles = 3.775*[12+4*35] = 573.8

volume required = 573.8/1.5 =382.33 litres

 

i hope this clears :)

krishnan s
13 years ago

thanks for the answer. but i did not understand why we multiply 4 with avogadro number